i am trying to find the y coordinate of the vertex. the x is -2 and when i plug it in to the equation i get
3(-2)^2+12(-2)+8 =-28, but the answer is supposed to be -4 can someone tell me what i am doing wrong?

Answer by rfer(12730)

I think you lost track of a sign some where.
3(-2)^2+12(-2)+8=
3(+4)-24+8=
12-24+8=
20-24=-4